OpenStudy (anonymous):

log(2x+1)-log(x-2)=1

OpenStudy (anonymous):

that would give log(2x+1/x-2)=1 becoming (2x+1)/(x-2)=0 ->2x+1=0,x\=2 -> x=-1/2

OpenStudy (anonymous):

\[\log(2+x)-\log(x-2)=1\\ \log\frac{2+x}{x-2}=1\\ \frac{2+x}{x-2}=10^1\\ 2+x=10x-20\\ 22=9x\\\\ x=\frac{22}{9} \]

OpenStudy (anonymous):

nope, 10^1=10 what you want is 10^0=1

OpenStudy (jiteshmeghwal9):

REMEMBER:\[\log_{e}{b}-\log_{e}{a}=\log_{e}\frac{b}{a} \]

OpenStudy (anonymous):

ohhh, boyyy !